OpenStudy (anonymous):

SO IF YOU GUYS ARE WILLING TO CLARIFY MY ANSWERS FOR THESE 4 QUESTIONS IT WOULD BE GREATLY BLESSED! btw i have my answers for them already i just need to make sure if they are right.

OpenStudy (anonymous):

the answer are the following: D,E,C,E

OpenStudy (anonymous):

@satellite73 @KingGeorge @joemath314159 @wolf1728 @radar @Luis_Rivera @genius12

OpenStudy (anonymous):

@mathcalculus @mathslover @Oreoo @o_nanners @souvik @harshadh

OpenStudy (anonymous):

btw i did them on my own! :)

OpenStudy (anonymous):

@nincompoop @nikakuz @wild @Jamierox4ev3r @agent0smith @ash2326 @calculusxy @codeblue11 @Luigi0210

OpenStudy (anonymous):

30 = d

OpenStudy (bambimonster):

28.e

OpenStudy (anonymous):

@Ashleyisakitty @souvik @shamim @souvik @thecollinjc1 @TotesMcGoats @ilfy214

OpenStudy (anonymous):

@bambimonster FOR 28 I WAS BETWEEN E & A 0_o they both give me 65 as the answer for the area

OpenStudy (anonymous):

28) E (x+3)(x-5) = 65 x^2-2 x-80 = 0 x = -8 , x = 10 Then check to see which one fits in the equation

OpenStudy (bambimonster):

@Big_Booty it should be e. because if u solve out x^2-2x-80=0 , u get x=10

OpenStudy (anonymous):

28) (10+3)(10-5) = 65

OpenStudy (anonymous):

Got it guys thanks! :)

OpenStudy (anonymous):

can you guys help me clarify 27 & 29?

OpenStudy (bambimonster):

4 servings:9beets 37servings:83.25beets so 1 borscht has 20, how many u need to add up to 83.25 ... im not really sure for this..

OpenStudy (anonymous):

for 27 i got E! i broke it down to serving=butchels & try to figure out how many bags i would need

OpenStudy (bambimonster):

how did u get e

OpenStudy (anonymous):

so i need 37 servings so i added 4 until i get to 37 & =to 9 then after that i calculated how many bags i would need

OpenStudy (anonymous):

sorry im confusing :(

OpenStudy (bambimonster):

hmm but i think u should do the ratio way to get the amount of beets u need .. like what i wrote above..

OpenStudy (bambimonster):

so it looks like u need 83.25 beets.. and 1 borscht has 20.. so u need about 4 borschtx 20 = 80.

OpenStudy (radar):

I never could see the problems.......where were they posted?

OpenStudy (anonymous):

@radar I'm guessing that she deleted the link?

OpenStudy (radar):

That must of been the case @ilfy214 as I never saw a link. Thanks.